### v1

# A Short Note on Upper Bounds for Graph Neural Operator Convergence Rates

Abstract
Graphons, as limits of graph sequences, provide a framework for analyzing the asymptotic behavior of graph neural operators. Spectral convergence of sampled graphs to graphons yields operator-level convergence rates, enabling transferability analyses of GNNs. This note summarizes known bounds under no assumptions, global Lipschitz continuity, and piecewise-Lipschitz continuity, highlighting tradeoffs between assumptions and rates, and illustrating their empirical tightness on synthetic and real data.
Index Terms—  graph neural operator, graphon, convergence rates, graph neural networks, transferability
1 Introduction
Graph neural networks (GNNs) are widely used in drug discovery [1, 2], social networks [3, 4], recommendation systems [5], and NLP [6, 7, 8]. GNNs operate on graph-structured data via message passing and aggregation [9], but training on large graphs is computationally expensive. Studying GNN behavior on families of large graphs, exploiting low-dimensional structure like finite rank or bandlimitedness, offers statistical guarantees and mitigates scalability challenges [10, 11].
A framework for analyzing families of large graphs is given by graphons: symmetric measurable functions ${\mathbf{W}}:[0,1]^{2}\to[0,1]$ representing the probability of an edge between two nodes placed in the interval $[0,1]$ [12, 13]. Graphons are limit objects of graph sequences under the cut distance, a metric that accounts for arbitrary relabelings, providing an avenue for analyzing the asymptotic behavior of graph neural operators and their transferability properties [14]. As graphons define bounded, symmetric operators, their spectra capture the frequencies on which graph neural operators act, and spectral convergence of graph sequences to graphon limits translates to operator convergence [15].
The spectral argument for operator convergence of GNNs on graph sequences uses two facts: (i) Weyl’s inequality bounds eigenvalue perturbations by operator-norm kernel differences [16]; and (ii) graphon convergence in cut distance implies convergence in cut and operator norms under relabeling [17, Proposition 4]. Together, these yield spectral convergence rates from cut distance, at a $\mathcal{O}(1/\sqrt[4]{\log n})$ rate [15, 18].
Sharper rates follow from structural assumptions on the graphon. Ruiz et al. [14] assume global Lipschitz continuity under fixed labeling, yielding $\mathcal{O}(\sqrt{\log n/n})$ rates, though this fails under arbitrary relabelings [12, Figure 4]. Avella Medina et al. [19] assume piecewise-Lipschitz structure, giving intermediate $\mathcal{O}(\sqrt[4]{\log n/n})$ rates while permitting flexible labeling within pieces. Lipschitz constants can be estimated in practice via sort-and-smooth or related methods [20, 21, 22].
In this short note, we collect and state these operator convergence rates in a unified framework. Our focus is on graph neural operator convergence, understood as eigenvalue stability under graphon limits. Section 2 introduces basic definitions. Section 3 presents the convergence rates under no assumptions, Lipschitz continuity, and piecewise-Lipschitz continuity. A comparison of the bounds is given in Section 3.4, with numerical illustrations in Section 4, and concluding remarks in Section 5.
2 Preliminary Definitions
We consider undirected graphs $G=(V,E,W)$ consisting of a set of nodes $V$, a set of edges $E\subseteq V\times V$, and a weight function $W:E\to\mathbb{R}$, assigning a real-valued weight $w_{ij}$ to each edge $(i,j)\in E$. An undirected graph $G=(V,E,W)$ can be represented by its symmetric adjacency matrix $A\in\mathbb{R}^{n\times n}$, where $A_{ij}=w_{ij}$ if $(i,j)\in E$, and $A_{ij}=0$ otherwise. We denote a sequence of graphs $\{G_{n}\}$, with $n\in{\mathbb{N}}\setminus\{0\}$.
2.1 Graphons
Graphons can be seen both as generative models for graphs and as limit objects of convergent graph sequences [15, 23]. Using a graphon as a generative model, we can construct an $n$-node graph $G_{n}$ from $\mathbf{W}$} by sampling points $u_{i}\in[0,1]$—associated with nodes $i\in[1,2,...,n]$—and assigning edge weight ${\mathbf{W}}(u_{i},u_{j})$ to edge $(i,j)$ to obtain a complete weighted graph; or, connecting edge $(i,j)$ with probability $p_{ij}={\mathbf{W}}(u_{i},u_{j})$ to obtain a stochastic unweighted graph. In either case, we say $G_{n}$ is sampled from $\mathbf{W}$}.
A sequence of graphs $\{G_{n}\}$ is said to converge if, for every fixed finite motif $F$ (such as a triangle or $k$-cycle), the proportion of adjacency-preserving mappings (homomorphisms) from $F$ into $G_{n}$ stabilizes as $n\to\infty$ [23, 15]. This proportion, called the homomorphism density $t(F,G_{n})$, measures how frequently $F$ appears in $G_{n}$. Graphons are the limits of such sequences, in that the densities of homomorphisms $t(F,G_{n})$ converge to the graphon homomorphism density. Explicitly, the graphon homomorphism density $t(F,\mathbf{W})$ represents the probability of sampling $F$ from the graphon $\mathbf{W}$. We say that $\{G_{n}\}$ converges to $\mathbf{W}$ if and only if $t(F,G_{n})\to t(F,\mathbf{W})$ for all $F$; in this case, $\mathbf{W}$ is the limit graphon.
2.2 Cut norm and cut distance
The cut norm of a graphon $\mathbf{W}$ measures the strongest concentration of connections between vertex subsets
$\|\mathbf{W}\|_{\Box}=\sup_{S,T\subset[0,1]}\left|\int_{S\times T}\mathbf{W}(u,v)\text{d}u\text{d}v\right|.$
The cut distance between two graphons, $\mathbf{U}$, $\mathbf{V}$, measures how different they are up to relabeling. It is defined as the infimum of their cut norm difference over all measure-preserving bijections $\phi:[0,1]\to[0,1]$:
$\delta_{\Box}(\mathbf{U},\mathbf{V})=\inf_{\phi\in\Phi}\|\mathbf{U}-\mathbf{V}^{\phi}\|_{\Box}.$ (1)
This accounts for arbitrary vertex labels and ensures the distance is label-invariant [15].
It can be shown that convergent graph sequences in the homomorphism density sense also converge in the cut distance. This statement is made formal by defining the graphons ${\mathbf{W}}_{G_{n}}$ induced by the graphs $G_{n}$, which allows comparison of the graphs $G_{n}$ and the graphon $\mathbf{W}$} through the above-defined distance. For an $n$-node graph $G$, such induced graphon is created by constructing a regular partition $I_{1}\cup...\cup I_{n}$ of $[0,1]$, where $I_{j}=[(j-1)/n,j/n],1\leq j\leq n-1$ and $I_{n}=[(n-1)/n,1]$, and defining
${\mathbf{W}}_{G}(u,v)=\sum_{j=1}^{n}\sum_{k=1}^{n}[A]_{jk}\mathbb{I}(u\in I_{j})\mathbb{I}(v\in I_{k}).$
If $\{G_{n}\}$ converges to $\mathbf{W}$} in the homomorphism density sense, $\delta_{\Box}({\mathbf{W}}_{G_{n}},{\mathbf{W}})\to 0\text{ as }n\to\infty$. This establishes the cut distance as the standard metric for graphon convergence [15, 23].
2.3 The graphon operator and Weyl’s inequality
Fig. 1: Piecewise-Lipschitz graphon with 200 nodes, $K=4$, and per-piece Lipschitz constant ${\mathcal{L}}_{k}\leq 4$.
As $\mathbf{W}$} is bounded and symmetric, it defines an integral linear operator $T_{{\mathbf{W}}}$ on $L^{2}([0,1])$. More specifically, given a function $X\in L_{2}([0,1])$—mapping “graphon nodes” $u\in[0,1]$ to the reals—the graphon operator acts on it as:
$T_{\mathbf{W}}X=\int_{0}^{1}{\mathbf{W}}(u,\cdot)X(u)\text{d}u.$ (2)
This is the operator that parametrizes graphon convolutions in graphon neural operators such as graphon neural networks (WNNs), which are the idealized limit objects of GNNs on sequences of graphs converging to graphons [14, 17, 18].
Since graphon operators are compact, self-adjoint operators on $L^{2}([0,1])$, their spectra consist of real eigenvalues that capture the frequencies on which graphon neural operators act [17]. To analyze convergence of graph neural operators to their graphon counterparts, we need control over how the spectra of sampled graph operators deviate from the limiting graphon operator. Weyl’s inequality adapted to graphons provides this type of stability guarantee, bounding the difference between eigenvalues of two operators by their operator-norm distance [16]. For a graph $G_{n}$, sampled from a graphon $\mathbf{W}$}, with induced graphon ${\mathbf{W}}_{G_{n}}$, we have
$|\lambda_{i}({\mathbf{W}})-\lambda_{i}({\mathbf{W}}_{G_{n}})|\leq\|T_{{\mathbf{W}}}-T_{{\mathbf{W}}_{G_{n}}}\|_{2}$ (3)
where in a slight abuse of notation we use $\|\cdot\|_{2}$ to denote the operator norm induced by the $L_{2}$ norm.
3 Graph Neural Operator Convergence Rates
This section presents operator-level convergence results for graph sequences to their graphon limits. The rates are derived by combining graphon convergence bounds with Weyl’s inequality (3), transferring these bounds to operator spectra. We focus on graph neural operator convergence, measured via eigenvalue stability. Each subsection states structural assumptions on the graphon and the resulting convergence rate.
3.1 Standard case
Without assumptions on the limit graphon $\mathbf{W}$}, GNN convergence bounds via Weyl’s inequality (3) follow directly from cut distance convergence, which can be related to the operator norm as summarized in Lemma 1. For a symmetric kernel ${\mathbf{K}}:[0,1]^{2}\to[-1,1]$ (e.g., ${\mathbf{W}}-{\mathbf{W}}^{\prime}$), this lemma connects its $L_{2}$-induced operator norm to its cut norm.
Lemma 1 (Adapted from [17]).
Let ${\mathbf{K}}:[0,1]^{2}\to[-1,1]$. Then,
$\|{\mathbf{K}}\|_{\Box}\leq\|T_{\mathbf{K}}\|_{2}\leq\sqrt{8\|{\mathbf{K}}\|_{\Box}}.$ (4)
For sequences of stochastic graphs $\{G_{n}\}$ sampled from $\mathbf{W}$} as described in Sec. 2.1, explicit convergence rates for the cut distance (1) are given by [15, Second Sampling Lemma], adapted in Lemma 2 below.
Lemma 2 (Second Sampling Lemma).
Let $G_{n}$ be a graph randomly sampled from an arbitrary graphon $\mathbf{W}$} with associated induced graphon ${\mathbf{W}}_{G_{n}}$. With probability at least $1-\exp(-n/(2\log n))$,
$\delta_{\square}({\mathbf{W}}_{G_{n}},{\mathbf{W}})=\inf_{\phi\in\Phi}\|{{\mathbf{W}}_{G_{n}}}-{\mathbf{W}}^{\phi}\|_{\Box}\leq\frac{22}{\sqrt{\log n}}$ (5)
where $\Phi$ is the set of measure-preserving bijections on $[0,1]$.
Combining Weyl’s inequality with Lemmas 1 and 2 recovers the first convergence rate for graphon neural operators, without any structural assumptions on $\mathbf{W}$}.
Proposition 3 (Standard case).
Let $G_{n}$ be a graph randomly sampled from a graphon $\mathbf{W}$} with induced graphon ${\mathbf{W}}_{G_{n}}$. With probability at least $1-\exp(-n/(2\log n))$, the eigenvalues of the corresponding operators satisfy
$|\lambda_{i}({\mathbf{W}}_{G_{n}})-\lambda_{i}({\mathbf{W}})|=\mathcal{O}\!\left(\frac{1}{\sqrt[4]{\log n}}\right).$
Proof.
Let $\phi$ be the measure preserving bijection achieving the infimum in Lemma 2. As applying any measure preserving bijection over the unit interval will not affect the spectrum of $T_{\mathbf{W}}$, by Weyl’s inequality (3) and Lemma 1 we have
[…]
Fig. 2: Examples of graphons. Left: Synthetic graphon $\textbf{f}=x\cdot y$ with Lipschitz constant $0.0265$ and largest piecewise-Lipschitz constant $0.353$. Center: Cora graphon with Lipschitz constant $60.653$ and largest piecewise-Lipschitz constant $99.5$. Right: PubMed graphon with Lipschitz constant $58.534$ and largest piecewise-Lipschitz constant $199.24$.
3.2 Lipschitz case
At the cost of generality (and, more importantly, exchangeability of nodes, as we discuss later in this section), we can impose Lipschitz continuity as an additional structural assumption on the graphon $\mathbf{W}$}. Under this assumption, the Hilbert-Schmidt (HS) norm difference between a graphon and a graph randomly sampled from it can be bounded as follows.
Lemma 4 (Adapted from [14]).
Let $G_{n}$ be a graph randomly sampled from a Lipschitz graphon $\mathbf{W}$} with associated induced graphon ${\mathbf{W}}_{G_{n}}$. With probability at least $1-\chi\times[1-2x_{1}]\times[1-x_{2}]$,
$\begin{split}\|{\mathbf{W}}-{\mathbf{W}}_{G_{n}}\|_{2}&\leq\frac{2{\mathcal{L}}_{{\mathbf{W}}}}{n}\log\!\left(\frac{(n+1)^{2}}{\log(1-x_{1})^{-1}}\right)\\ &\quad+\frac{1}{n}\sqrt{4n\log\!\Big(\tfrac{2n}{\chi}\Big)},\end{split}$ (7)
where ${\mathcal{L}}_{{\mathbf{W}}}$ is the Lipschitz constant of the graphon, $x_{1},x_{2}\in(0,0.3]$, $\chi>0$, and $n\geq\tfrac{4}{x_{2}}$.
Combining this bound with Weyl’s inequality (3) yields the following operator convergence rate.
Proposition 5 (Lipschitz case).
Let $G_{n}$ be a graph randomly sampled from a Lipschitz graphon $\mathbf{W}$} with induced graphon ${\mathbf{W}}_{G_{n}}$. With high probability,
$|\lambda_{i}({\mathbf{W}}_{G_{n}})-\lambda_{i}({\mathbf{W}})|=\mathcal{O}\!\left(\sqrt{\tfrac{\log n}{n}}\right).$
Proof.
Weyl’s inequality gives $|\lambda_{i}({\mathbf{W}}_{G_{n}})-\lambda_{i}({\mathbf{W}})|\leq\|T_{{\mathbf{W}}_{G_{n}}}-T_{{\mathbf{W}}}\|_{2}$. Since $\|T_{\mathbf{K}}\|_{2}\leq\|{\mathbf{K}}\|_{2}$, Lemma 4 directly implies $\|T_{{\mathbf{W}}_{G_{n}}}-T_{{\mathbf{W}}}\|_{2}\leq\|{\mathbf{W}}_{G_{n}}-{\mathbf{W}}\|_{2}$, which yields the stated rate $\mathcal{O}(\sqrt{\log(n)/n})$. ∎
3.3 Piecewise-Lipschitz case
A more forgiving assumption is piecewise-Lipschitz continuity. As in [19], we consider piecewise-Lipschitz graphons $\mathbf{W}$} that can be partitioned into at most $K^{2}$ “pieces,” where $K$ is the total number of non-overlapping intervals in the partition. Each piece is Lipschitz with constant ${\mathcal{L}}_{k}$, and we let ${\mathcal{L}}=\max_{k}{\mathcal{L}}_{k}$ for consistency with the global Lipschitz notation in Sec. 3.2. An example is presented in Figure 1.
Note that the different pieces of the graphon can have different Lipschitz constants, and $\mathbf{W}$} can be discontinuous along partition boundaries. Under the piecewise-Lipschitz assumption, the HS norm difference between a graphon and a graph randomly sampled from it can be bounded as follows.
Lemma 6 (Adapted from [17, 19]).
Let $G_{n}$ be a graph randomly sampled from a piecewise-Lipschitz graphon $\mathbf{W}$} with associated induced graphon ${\mathbf{W}}_{G_{n}}$. With probability at least $(1-\chi)(1-\delta^{\prime})$,
$\begin{split}\|{\mathbf{W}}-{\mathbf{W}}_{G_{n}}\|_{2}&\leq 2\sqrt{({\mathcal{L}}^{2}-K^{2}){d_{n}}^{2}+Kd_{n}}\\ &\quad+\frac{1}{n}\sqrt{4n\log\!\big(\tfrac{2n}{\chi}\big)},\end{split}$ (8)
where $\delta^{\prime}\in(ne^{-n/5},e^{-1})$, $\chi>0$, and $d_{n}=\tfrac{1}{n}+\sqrt{\tfrac{8\log(n/\delta)}{n+1}}$.
Combining this bound with Weyl’s inequality (3) yields the following operator convergence rate.
Proposition 7 (Piecewise-Lipschitz case).
Let $G_{n}$ be a graph randomly sampled from a piecewise-Lipschitz graphon $\mathbf{W}$} with induced graphon ${\mathbf{W}}_{G_{n}}$. With high probability,
$|\lambda_{i}({\mathbf{W}}_{G_{n}})-\lambda_{i}({\mathbf{W}})|=\mathcal{O}\!\left(\sqrt[4]{\tfrac{\log n}{n}}\right).$
Proof.
Weyl’s inequality gives $|\lambda_{i}({\mathbf{W}}_{G_{n}})-\lambda_{i}({\mathbf{W}})|\leq\|T_{{\mathbf{W}}_{G_{n}}}-T_{{\mathbf{W}}}\|_{2}$. Since $\|T_{\mathbf{K}}\|_{2}\leq\|{\mathbf{K}}\|_{2}$, Lemma 6 implies the stated rate $\mathcal{O}(\sqrt[4]{\tfrac{\log n}{n}})$. ∎
3.4 Comparison and discussion
The bounds presented in Section 3 illustrate the impact of imposing additional structural assumptions on the graphon $\mathbf{W}$}, particularly as $n$ grows. Imposing structural assumptions results in substantially faster bounds than those obtained for arbitrary graphons. In particular, the bounds shrink faster under the piecewise-Lipschitz assumption, and even more rapidly under the global Lipschitz assumption.
The role of permutations, or node relabelings, also differs across these settings. Under standard assumptions, the bounds are permutation-invariant: reordering the inputs has no effect. Under the Lipschitz assumption, however, permutation invariance is lost; a suitable labeling of inputs must either be provided or learned. In the piecewise-Lipschitz case, the bounds are invariant to permutations between pieces, but not to permutations within each piece.
GNN transferability bounds typically decompose into two factors: a constant term that depends on structural and architectural assumptions (such as GNN depth), and a rate term that decays asymptotically with $n$, determined by spectral convergence. The focus of this note is on the operator-level rates, which can be directly substituted into any existing transferability bound, e.g., [14], to yield guarantees under different structural assumptions on the underlying graphon.
When considering the application of the operator to a specific input signal (i.e., the function $X$ in (2)), a further consideration in transferability is input signal convergence. In the fixed labeling case, this has been established by showing that graph signals converge to Lipschitz graphon signals in $L_{2}$ [14, 17]. In the general case, Levie et al. [18] introduced a cut metric for graphon–signal pairs, which controls convergence of both structure and signals simultaneously. Importantly, incorporating these signal convergence errors does not alter the rates derived here, with the resulting transferability bounds inheriting exactly the same asymptotic decay rates.
4 Numerical examples
We illustrate bound performance using three examples: a synthetic graphon, the Cora dataset, and the PubMed dataset. The synthetic graphon is generated from a smooth function over $0,1]^{2}$ to define connection probabilities. For Cora, we use the full adjacency matrix as the graphon, while for PubMed we use the adjacency matrix of a randomly sampled subset. We can normalize the graphon adjacency matrices, via the $L_{1}$ norm, and reorder by node degree. Leveraging polynomial interpolation we determine an approximate Lipschitz constant for the graphons and particular graphon “pieces”. Sorting and interpolating is consistent with the sort-and-smooth approach used in graphon estimation literature [20].
Figure 2 compares the first three eigenvalues of subgraphs sampled from the graphons to the graphon eigenvalues and the upper bounds. Bounds with additional assumptions perform better, and although the standard bound initially appears tighter than the piecewise-Lipschitz bound, the latter eventually overtakes it. None of the bounds are tight, indicating that incorporating more graph structure could improve them.
5 Conclusion
Upper bounds on GNN convergence rates are key for analyzing convolutional processing on graph sequences with graphon limits. Few approaches exist, with major contributions by [14], [18], and [19] presented in the last decade. In this work, we present and compare different graphon-based bounds both theoretically and via numerical examples. These bounds extend to large-scale graphs, offering new opportunities to improve GNN convergence and transferability in real-world tasks.